### v1

@@ 3.3.2 Constructing input projectors
The role of input projectors is to project data from different modalities into a common feature space. As discussed in section 2.3, modality alignment can be achieved through either fusion encoders or dual encoders. When constructing input projectors, the key decision is whether to use a bridging strategy to integrate input vectors from different modalities or to use a fine-tuning approach to bring the projectors of different modalities closer together. These two strategies correspond to the concepts of fusion encoders and dual encoders, respectively. During training, using loss functions from multi-modality understanding tasks, such as multi-modality classification or generation task losses, can train the model’s cross-modality projection capability. Additionally, an end-to-end training approach can optimize the overall performance of the foundation model and train cross-modality projections at the same time. As previously mentioned, the CoDi-2 model [21] utilizes ImageBind [25] encoders aligned by CLIP as the image and audio modality encoders and part of the input projector. It then incorporates an MLP as another part of the input projector. During the end-to-end training process of the foundation model, the MLP is optimized, so as to align the image and audio to text.
3.3.3 Constructing the backbone calculator
@@
The backbone calculator is responsible for understanding and generating feature vectors of the central modality. Constructing a backbone calculator for a specific domain begins with identify the most common and information-rich data modality in that domain as the processed modality by the backbone calculator. The model architecture is then designed based on this modality. Currently, mainstream model architectures are based on Transformers. A complete Transformer model consists of an encoder and a decoder, where the encoder analyzes the input data to extract compact feature representations, and the decoder uses these feature representations to generate the output content. Since the structures of the encoder and decoder are different, generally, the encoder has stronger comprehension capabilities, while the decoder has more powerful generative capabilities. Transformer-based foundation model backbone calculators have three main architectural forms, including encoder-only architectures, decoder-only architectures, and encoder-decoder architectures. The characteristics of these three architectures are summarized in Table 5.
Table 5: Comparison between encoder-only structure, decoder-only structure and encoder-decoder structure
@@
    Backbone processor based on encoder-only architecture: The encoder-only architecture consists solely of the encoder part of the Transformer. It is typically used for tasks that require understanding input text rather than generating new text sequences, such as text classification and sentiment analysis. Due to containing only the encoder part, the encoder model structure is relatively simple but can only produce fixed-length outputs, resulting in limited generation capability. In terms of generation tasks, backbone processors based on encoder-only model can only handle tasks such as sequence completion. BERT [7] is a well-known example of an encoder-only model.
@@
    Backbone processor based on decoder-only architecture: In the decoder-only architecture, the decoder directly processes the input sequence and generates the output sequence without a separate encoder to compactly represent the input sequence. This reduces parameter count and computational overhead but also makes it more challenging for the model to understand input sequences, thus limiting its ability to handle long sequences. This architecture does not require explicit context representation when generating output sequences but captures information automatically within the sequence through self-attention mechanisms. Backbone processors based on decoder-only architecture typically employ autoregressive generation, meaning they generate words or characters one by one based on previous generated content to complete sequence text generation tasks. Models like the GPT series [4, 5, 6, 22] and the LLaMA series [10, 11] of foundation language models belong to the decoder-only architecture.
@@
    Backbone processor based on encoder-decoder architecture: The encoder-decoder architecture can simultaneously possess the understanding capability of an encoder and the generation capability of a decoder, but also results in higher parameter count and computational costs. Models like Meta’s BART [15] and Google’s T5 [16] adopt this architecture.
3.3.4 Constructing output projectors and modality decoders
Modality decoders come in two types: generative and discriminative. Generative modality decoders can produce high-quality data samples given conditional information. Discriminative modality decoders, on the other hand, excel in precise reconstruction of data samples based on input vectors. The type of modality decoders also affects the design of output projectors. Thus it requires a joint consideration of these two modules. Figure 7 (a) and (b) respectively illustrate the operational processes of generative and discriminative modality decoders.
Figure 7: The workflow of output projector and modality decoder. (a) Generative modality decoder;(b) Discriminative modality decoder.
@@ 3.3.4 Constructing output projectors and modality decoders
    Generative modality decoder: A generative modality decoder utilizes a generative model as its neural network, which can control the generation process using conditional information. It takes other modality features as conditional information and generated data as decoding output. Such a generative model can be termed as a generative modality decoder, without need of explicit construction of output projectors. For instance, DiT [50], a diffusion-based image generation model, can progressively generate images based on previous results and conditional vectors, where cross-attention mechanisms serve the function of the output projector. Similarly, the VAR model [89] also uses self-attention mechanism as the output projector function. By incorporating modality labels in the starting position, VAR knows which content needs to be controlled for generation and generate images in a autoregressive way.
    Generative modality decoders typically employ end-to-end training strategies. During training, the generation part of the model and the modality interaction part are optimized simultaneously. The training objective usually involves minimizing the difference between generated data and real data while ensuring that generated data meets given conditions. For example, if the model’s objective is to generate images based on textual descriptions, a large number of text-image pairs are used during training. And the model is optimized by comparing the similarity between generated images and real images. This similarity can be measured using pixel-level loss functions (such as mean squared error) or more advanced perceptual losses (such as VGG loss). Additionally, adversarial training can be employed to enhance generation quality.
@@
    Discriminative modality decoder: A discriminative modality decoder is responsible only for directly reconstructing feature vectors into their original data form. Thus, it requires an explicit output projector to convert feature vectors from other modalities to the target modality for use. For example, when using the decoder of VQGAN [48] as the discriminative modality decoder in a multi-modality foundation model, it is necessary to first explicitly construct an output projector to transform feature vectors from other modality domains to the image modality, and then use the decoder portion to decode the feature vectors into their original data form. This explicit output projector is typically trained using supervised way, where the model takes feature vectors from other modalities as input and learns to project these features into feature vectors of the target modality by minimizing the error between the two feature vector. Besides, modality decoders are trained together with modality encoders as autoencoders, aiming to improve reconstruction performance by minimizing reconstruction errors.
@@ 4 Applications of domain-specific foundation models
Foundation models have emerged as powerful tools with extensive application prospects across various domains. These foundation models possess the capability not only to handle massive data and complex tasks but also to bring about new breakthroughs and innovations.
@@
  • Telecommunications: Foundation models are expected to be widely applied in sensing, transmission and network performance optimization, etc. For example, LLM fine-tuned for the telecommunications domain can be used to process network log data, model and solve specific network problems [44]. Additionally, foundation models in telecommunications, through the utilization of spatio-temporal correlations and knowledge reasoning [45], are expected to identify and prevent experience issues caused by degraded service quality, and delayed fault response times. This lays the foundation for fine-grained network optimization in real-time. In industrial scenarios, the business understanding capability of foundation models is also expected to help optimize signal transmission and scheduling strategies, improving network efficiency. In research on network foundation models, the NetGPT architecture proposed in the article [90] is expected to become an effective approach to achieve endogenous intelligence in telecommunications networks, while the article [91] discusses the challenges and issues that may be encountered in constructing foundation models for telecommunications. The NetLLM proposed in the article [92] adapts LLM to serve several specific downstream communication tasks. [93] introduces a new memory mechanism to facilitate the embedding of foundation models into the semantic communication process. By leveraging the foundation models’ capability to understand and generate multimodal data, it becomes feasible to propose new and more complex transmission approaches that enhance transmission performance.
@@
  • Autonomous driving: Foundation models play a core role in various key aspects such as vehicle perception, decision-making, and motion control [46]. Specifically, perception tasks in autonomous driving involve real-time monitoring of the vehicle’s surroundings, including other vehicles, pedestrians, traffic signs, and road conditions. Foundation models, by analyzing data collected from sensors such as cameras, radar, and LiDAR (Light Detection and Ranging), can identify various objects and construct detailed maps of the vehicle’s surroundings. This advanced perception capability is fundamental to achieving safe autonomous driving. At the decision-making level, foundation models need to make rapid and accurate judgments based on perceived information, such as avoiding obstacles, selecting appropriate driving paths, and devising optimal driving strategies in complex traffic situations. Multi-modality foundation models like DriveGPT [94] can not only process visual data but also understand and respond to language-mode instructions, such as planning routes based on voice-input destinations. Additionally, pFedLVM [95] can utilize the powerful performance of pre-trained visual foundation models for image feature extraction, serving as the basis for down-stream tasks.
@@
  • Mathematics: Since many general-purpose LLMs are typically pre-trained on datasets that include open-source mathematical corpora, they inherently possess some capacity to address mathematical problems, and through domain-specific enhancements, this capacity can be significantly improved [38, 96, 97, 98, 99]. The MAmmoTH proposed in [97] combined chain-of-thought and program-of-thought, fully leveraging the understanding capability of large language models and the computational power of programming languages, achieving good performance in mathematical reasoning. [98] indicates through experiments that the seamless integration of LLMs’ programming and reasoning abilities enables them to progressively model and solve complex problems.
@@
  • Medicine: The applications of foundation models in medicine encompass various aspects, including disease diagnosis, patient treatment, analysis and prediction of genetic and protein structure data, and medical education [40, 100, 101, 102, 103, 41]. The HuatuoGPT proposed in the article [102] can simulate the diagnosis and treatment process of doctors, providing preliminary medical consultation and advice for patients. This model not only reduces the workload of doctors but also enables patients to receive timely medical services in remote or resource-limited environments. [103] proposes BiomedGPT, which is capable of performing multimodal medical tasks and has achieved excellent performance in understanding and generating medical-related content.
@@
  • Law: Foundation models can conduct in-depth analysis of legal documents, identify key information and legal concepts in the text, thereby assisting lawyers and legal advisors in more precise case analysis and legal consultation [37, 104, 105]. For example, foundation models can identify clauses in contracts, extract important legal elements such as obligations, rights and conditions, help lawyers quickly understand document contents, and identify potential legal risks. Moreover, foundation models can be used for case logical reasoning, predicting possible outcomes of cases by analyzing historical cases and relevant legal provisions, providing data support for lawyers to formulate defense strategies. The ChatLaw [106] can provide real-time legal consultation and answers, helping non-professionals understand complex legal issues and even generate drafts of legal documents, reducing the workload of lawyers. Additionally, foundation models can assist in legal research, quickly retrieve relevant legal literature and precedents, and provide solid evidence for legal arguments.
@@
  • Arts: The application of foundation models is exploring and changing the ways of creative expression and the process of artistic production. By learning from a large number of artworks and creative concepts, foundation models can generate novel artworks, music, literary works, etc., providing creative inspiration and support for artists [42]. For example, generative models can be used to produce artworks [107, 108], and fine-tuning techniques can be applied to adjust the style and content of the generated works [109, 110]. Currently, in the field of video generation, Sora developed by OpenAI allows users to control generated content using text, producing lifelike video works. Furthermore, foundation models also show promising performance in art understanding [111, 112], which provides great potential in the field of art teaching and study.
@@
  • Finance: Foundation models can cover various tasks such as risk modeling, investment strategy management, market forecasting, etc., providing powerful tools for financial institutions and investors to optimize decision-making [39, 113]. For example, using foundation models to build credit scoring systems to assess borrowers’ credit risks. These models analyze factors such as borrowers’ historical credit records, financial conditions, and debt levels to predict their ability to repay loans, and based on this, decide whether to approve loan applications and the loan interest rates. Alternatively, foundation models can be used to consider the historical performance, correlations, risks, and expected returns of various asset classes, as well as historical market data, macroeconomic indicators, political events, etc., to make optimal decisions for investors. BloombergGPT, as detailed in [114], was pre-trained on a large-scale financial dataset, resulting in superior performance across many benchmarks. [115] introduces FinGPT, an open-source financial LLM, and positions it not only as a model but also as an open-source framework for Financial LLMs (FinLLMs), which has the potential to fuel innovation among researchers.
5 Future directions

### v2

@@ 3.3.2 Constructing input projector
The role of IPs is to project data from different modalities into a common feature space. As discussed in section 2.2, modality alignment can be achieved through either fusion encoders or dual encoders. When constructing IPs, the key decision is whether to use a bridging strategy to integrate input vectors from different modalities or to use a fine-tuning approach to bring the projectors of different modalities closer together. These two strategies correspond to the concepts of fusion encoders and dual encoders, respectively. During training, using loss functions from multi-modality understanding tasks, such as multi-modality classification or generation task losses, can train the model’s cross-modality projection capability. Additionally, an end-to-end training approach can optimize the overall performance of the FM and train cross-modality projections at the same time. As previously mentioned, the CoDi-2 model [21] utilizes ImageBind [25] encoders aligned by CLIP as the image and audio MEs and part of the IP. It then incorporates an MLP as another part of the IP. During the end-to-end training process of the FM, the MLP is optimized to align the image and audio to text.
3.3.3 Constructing backbone calculator
@@
The BC is capable of understanding and generating feature vectors of the central modality. Constructing a BC for a specific domain begins with identifying the most common and information-rich data modality in that domain as the processed modality by the BC. The model architecture is then designed based on this modality. Currently, mainstream model architectures are based on Transformers. A complete Transformer model consists of an encoder and a decoder, where the encoder analyzes the input data to extract compact feature representations, and the decoder uses these feature representations to generate the output content. Since the structures of the encoder and decoder are different, generally, the encoder has stronger comprehension capabilities, while the decoder has more powerful generative capabilities. Transformer-based FM BCs have three main architectural forms: encoder-only architectures, decoder-only architectures, and encoder-decoder architectures. The characteristics of these three architectures are summarized in Table 6.
Table 6: Comparison between encoder-only structure, decoder-only structure, and encoder-decoder structure
@@
    BC based on encoder-only architecture: The encoder-only architecture consists solely of the encoder part of the Transformer. It is typically used for tasks that require understanding input text rather than generating new text sequences, such as text classification and sentiment analysis. Due to containing only the encoder part, the encoder model structure is relatively simple but can only produce fixed-length outputs, resulting in limited generation capability. Regarding generation tasks, BCs based on an encoder-only model can only handle tasks such as sequence completion. BERT [7] is a well-known example of an encoder-only model.
@@
    BC based on decoder-only architecture: In the decoder-only architecture, the decoder directly processes the input sequence and generates the output sequence without a separate encoder to compactly represent the input sequence. This reduces parameter count and computational overhead but also makes it more challenging for the model to understand input sequences, thus limiting its ability to handle long sequences. This architecture does not require explicit context representation when generating output sequences but captures information automatically within the sequence through self-attention mechanisms. BCs based on decoder-only architecture typically employ autoregressive generation, meaning they generate words or characters one by one based on previously generated content to complete sequence text generation tasks. Models like the GPT series [4, 5, 6, 22] and the LLaMA series [10, 11] of foundation language models belong to the decoder-only architecture.
@@
    BC based on encoder-decoder architecture: The encoder-decoder architecture can simultaneously possess the understanding capability of an encoder and the generation capability of a decoder, but also results in higher parameter count and computational costs. This architecture is adopted by models like Meta’s BART [15] and Google’s T5 [16].
3.3.4 Constructing output projector and modality decoder
MDs come in two types: generative and discriminative. Generative MDs can produce high-quality data samples given conditional information. Discriminative MDs, on the other hand, excel in the precise reconstruction of data samples based on input vectors. The type of MDs also affects the design of OPs. Thus it requires a joint consideration of these two modules. Figure 11 (a) and (b) respectively illustrate the operational processes of generative and discriminative MDs.
Figure 11: The workflow of OP and MD. (a) Generative MD;(b) Discriminative MD.
@@ 3.3.4 Constructing output projector and modality decoder
    Generative MD: A generative MD utilizes a generative model as its neural network, which can control the generation process using conditional information. It takes other modality features as conditional information and generates data as decoding output. Such a generative model can be termed as a generative MD without the need for explicit construction of OPs. For instance, DiT [51], a diffusion-based image generation model, can progressively generate images based on previous results and conditional vectors, where cross-attention mechanisms serve the function of the OP. Similarly, the VAR model [126] also uses the self-attention mechanism as the OP function. By incorporating modality labels in the starting position, VAR knows which content needs to be controlled for generation and generates images in an autoregressive way.
    Generative MDs typically employ end-to-end training strategies. During training, the generation part of the model and the modality interaction part are optimized simultaneously. The training objective usually involves minimizing the difference between generated data and real data while ensuring that generated data meets given conditions. For example, if the model’s objective is to generate images based on textual descriptions, a large number of text-image pairs are used during training. The model is optimized by comparing the similarity between generated images and real images. This similarity can be measured using pixel-level loss functions (such as mean squared error) or more advanced perceptual losses (such as VGG loss). Additionally, adversarial training can be employed to enhance generation quality.
@@
    Discriminative MD: A discriminative MD is only for directly reconstructing feature vectors into their original data form. Thus, it requires an explicit OP to convert feature vectors from other modalities to the target modality for use. For example, when using the decoder of VQGAN [127] as the discriminative MD in a multi-modality FM, it is necessary to first explicitly construct an OP to transform feature vectors from other modality domains to the image modality, and then use the decoder portion to decode the feature vectors into their original data form. This explicit OP is typically trained using a supervised way, where the model takes feature vectors from other modalities as input and learns to project these features into feature vectors of the target modality by minimizing the error between the two feature vectors. Besides, MDs are trained together with MEs as autoencoders, aiming to improve reconstruction performance by minimizing reconstruction errors.
@@ 4 Applications of domain-specific foundation models
FMs have emerged as powerful tools with extensive application prospects across various domains. These FMs possess the capability not only to handle massive data and complex tasks but also to bring about new breakthroughs and innovations.
@@
  • Telecommunications: FMs are expected to be widely applied in sensing, transmission, network performance optimization, etc. For example, an LLM fine-tuned for the telecommunications domain can be used to process network log data, model, and solve specific network problems [44, 128]. Additionally, FMs in telecommunications, through the utilization of spatio-temporal correlations and knowledge reasoning [45], are expected to identify and prevent experience issues caused by degraded service quality and delayed fault response times. Recent works also successfully incorporate telecommunication standards into LLMs’ reasoning [114, 129]. As a pioneering work on how to introduce knowledge of the wireless communication field to LLMs, [130] presents a comprehensive framework that elaborates on the principles and methods for constructing a “WirelessLLM” and provides examples to demonstrate its functionality. Specifically, TelecomGPT[131] acquires domain knowledge in telecom-specific datasets through continual pre-training, instruction tuning, and alignment tuning, after which it is able to answer related questions and generate codes in the telecom domain. This lays the foundation for fine-grained network optimization in real time. We illustrate the advantages of domain-specific enhancement by presenting the results from [131], as summarized in Table 7. This table highlights the performance of LLMs that have undergone domain-specific enhancements, denoted by TP (continual pretraining), TI (instruct tuning), and TA (alignment tuning) in their names. It is evident that these enhancements lead to significant improvements in performance across various metrics. The efficiency of foundation models is crucial in use [132]. In response to the significant challenges in data and communication faced when training FMs, [133] proposes multiple new paradigms to explore how Federated Learning can be utilized in a wireless context to address these challenges. In industrial scenarios, the business understanding capability of FMs is also expected to help optimize signal transmission and scheduling strategies, improving network efficiency. In research on network FMs, the NetGPT architecture proposed in the article [134] is expected to become an effective approach to achieve endogenous intelligence in telecommunications networks, while the article [135] discusses the challenges and issues that may be encountered in constructing FMs for telecommunications. The NetLLM proposed in the article [136] adapts LLM to serve several specific downstream communication tasks. [137] introduces a new memory mechanism to facilitate the embedding of FMs into the semantic communication process. By leveraging the FMs’ capability to understand and generate multimodal data, it becomes feasible to propose new and more complex transmission approaches that enhance transmission performance. In telecommunications, training domain-specific FMs on large datasets from telecommunication systems involves significant costs for data storage, processing power, and specialized hardware. Deployment costs are often high due to the need for real-time processing capabilities, especially in edge environments [138]. Additionally, ongoing operational costs arise from continuous monitoring, model updating, and adaptation to evolving network conditions.
    Table 7: Performance comparison of Telecom multiple-choice question (MCQ) answering on TeleQnA [114] dataset of mainstream LLMs and different telecom-specific LLMs obtained during different stages of TelecomGPT’s training pipeline
    LLMs Lexicon Research Overview Research Publications Standards Overview Standards Specifications Avg Llama3-8B 72 51.92 65.11 56.45 36.17 56.20 Llama3-8B-Instruct 80 67.31 69.77 59.68 50 64.80 Llama3-8B-TI 96 69.23 74.88 74.19 56.38 71.20 Llama3-8B-TI-TA 92 73.08 71.63 72.58 58.51 70.60 Mistral-7B 72 49.04 51.16 50 34.04 48.40 Mistral-7B-Instruct 84 64 65 56 51 62 Mistral-7B-TI 84 67.3 70.69 56.45 51.06 65.2 Mistral-7B-TI-TA 84 70.19 73.95 61.29 48.94 64 LlaMA-2-7B 62.5 52.24 49.18 48.28 40 48.94 LlaMA-2-7B-TI 84 57.69 63.26 56.45 50 59.80 LlaMA-2-7B-TP-TI 81.82 63.92 67 70 47.48 63.79
@@
  • Autonomous driving: FMs play a core role in various key aspects such as vehicle perception, decision-making, and motion control [46]. Specifically, perception tasks in autonomous driving involve real-time monitoring of the vehicle’s surroundings, including other vehicles, pedestrians, traffic signs, and road conditions [139]. FMs, by analyzing data collected from sensors such as cameras, radar, and LiDAR, can identify various objects and construct detailed maps of the vehicle’s surroundings. This advanced perception capability is fundamental to achieving safe autonomous driving. At the decision-making level, FMs need to make rapid and accurate judgments based on perceived information, such as avoiding obstacles, selecting appropriate driving paths, and devising optimal driving strategies in complex traffic situations [140, 141]. Multi-modality FMs like DriveGPT [142] can not only process visual data but also understand and respond to language-mode instructions, such as planning routes based on voice input destinations. Additionally, pFedLVM [143] can utilize the powerful performance of pre-trained visual FMs for image feature extraction, serving as the basis for downstream tasks. Considering the scarcity of high-quality public datasets for training LLMs in autonomous driving scenarios, [144] employs federated instruction tuning and expands the dataset by generating new instruction tracking data to mitigate the data scarcity. DriveMLM [139] is an LLM-based framework that can perform close-loop autonomous driving. It uses driving rules, user commands, and sensor inputs as its multi-modality LLM core inputs to model behavior planning. It standardizes decision states to connect language decisions with vehicle control commands. Training domain-specific FMs for autonomous driving requires extensive labeled data collection from real-world scenarios, which is costly and time-consuming. Additionally, the training demands substantial computational resources, particularly with large sensor data like LiDAR, radar, and camera feeds. Deploying these models involves significant expenses for hardware integration, safety certifications, and real-time vehicle processing.
@@
  • Mathematics: Since many general-purpose LLMs are typically pre-trained on datasets that include open-source mathematical corpora, they inherently possess some capacity to address mathematical problems, and through domain-specific enhancements, this capacity can be significantly improved [38, 145, 146, 147, 148]. The MAmmoTH proposed in [146] combined chain-of-thought and program-of-thought, fully leveraging the understanding capability of large language models and the computational power of programming languages, achieving good performance in mathematical reasoning. To provide a clear demonstration of the benefits of domain-specific enhancements, we collect the results from [146] in Table 8. MAmmoTH-Coder shows a clear performance advantage with general-purpose open-source LLMs and is comparable with closed-source LLMs. [147] indicates through experiments that the seamless integration of LLMs’ programming and reasoning abilities enables them to model and solve complex problems progressively. Due to the inherent complexity of mathematical problems and the need for vast, specialized datasets, training domain-specific foundation models in mathematics for tasks such as theorem proving or algorithm optimization can be challenging.
    Table 8: Experimental results on mathematical evaluation datasets including math problems from elementary, high school, and college levels. Math-SFT? means whether the model has been instruction-tuned on any math reasoning datasets
    Model Base Math-SFT? GSM8K MATH AQuA NumGLUE Avg Llama-1 - No 10.7 2.9 22.6 24.7 15.5 Llama-2 - No 14.6 2.5 30.3 29.9 19.3 Galactica-6.7B GAL GAL-Instruct 10.2 2.2 25.6 25.8 15.9 Code-Llama (PoT) - No 25.2 13.0 24.0 26.8 22.2 AQuA-SFT Llama-2 AQuA 11.2 3.6 35.6 12.2 15.6 Llama-1 RFT Llama-1 GSM8K 46.5 5.2 18.8 21.1 22.9 WizardMath Llama-2 GSM8K+MATH 54.9 10.7 26.3 36.1 32.0 MAmmoTH Llama-2 MathInstruct 53.6 31.5 44.5 61.2 47.7 MAmmoTH-Coder Code-Llama MathInstruct 59.4 33.4 47.2 66.4 51.6
@@
  • Medicine: The applications of FMs in medicine encompass various aspects, including disease diagnosis, patient treatment, analysis and prediction of genetic and protein structure data, and medical education [40, 149, 150, 151, 152, 41]. The HuatuoGPT proposed in the article [151] can simulate the diagnosis and treatment process of doctors, providing preliminary medical consultation and advice for patients. This model not only reduces the workload of doctors but also enables patients to receive timely medical services in remote or resource-limited environments. To illustrate the advantages of domain-specific enhancements effectively, we present part of the experimental results from [151] in Table 9. When contrasted with LLMs without medical domain-specific enhancements, HuatuoGPT demonstrates distinct comparative advantages. [152] proposes BiomedGPT, which is capable of performing multimodal medical tasks and has achieved excellent performance in understanding and generating medical-related content. In the medical area, the training datasets are often expensive to acquire due to the stringent requirements for privacy and regulatory compliance. Moreover, it is expensive to pay medical experts to annotate data.
    Table 9: Experimental results on Chinese medical QA dataset [153]
    Dataset Model BLEU-1 BLEU-2 BLEU-3 BLEU-4 GLEU ROUGE-1 ROUGE-2 ROUGE-L Distinct-1 Distinct-2 cMedQA2 T5 (fine-tuned) 20.88 11.87 7.69 5.09 7.62 27.16 9.30 20.11 0.41 0.52 DoctorGLM 13.51 7.10 3.72 2.00 5.11 22.78 5.68 12.22 0.85 0.96 ChatGPT 19.21 7.43 3.14 1.24 5.06 20.13 3.10 12.57 0.69 0.99 ChatGLM-6B 24.90 12.74 6.99 3.87 8.49 28.52 7.19 18.21 0.68 0.99 Ziya-LLaMA-13B 27.03 13.87 7.48 4.09 7.77 28.24 7.10 14.81 0.78 0.93 HuatuoGPT 27.39 14.38 8.06 4.55 8.52 29.26 8.02 15.46 0.74 0.93 webMedQA T5 (fine-tuned) 21.42 13.79 10.06 7.38 8.94 31.00 13.85 25.78 0.37 0.46 DoctorGLM 9.91 5.20 2.78 1.54 4.67 23.01 5.68 11.96 0.84 0.95 ChatGPT 18.06 6.74 2.73 1.09 4.71 20.01 2.81 12.58 0.65 0.87 ChatGLM-6B 23.42 12.10 6.73 3.83 8.04 28.30 6.87 18.49 0.63 0.87 Ziya-LLaMA-13B 22.16 11.70 6.53 3.74 6.91 27.41 6.80 13.52 0.76 0.93 HuatuoGPT 24.85 13.42 7.72 4.51 7.50 28.30 7.72 14.50 0.73 0.93 Huatuo-26M T5 (fine-tuned) 26.63 16.74 11.77 8.46 11.38 33.21 13.26 24.85 0.51 0.68 DoctorGLM 11.50 6.00 3.14 1.69 4.65 22.39 5.47 12.14 0.85 0.96 ChatGPT 18.44 6.95 2.87 1.13 4.87 19.60 2.82 12.46 0.69 0.89 ChatGLM-6B 24.46 12.75 7.20 4.13 8.50 28.44 7.31 18.58 0.67 0.89 Ziya-LLaMA-13B 25.58 13.39 7.46 4.24 7.30 28.14 7.18 14.78 0.77 0.93 HuatuoGPT 27.42 14.84 8.54 4.96 8.01 29.16 8.29 15.84 0.74 0.93
@@
  • Law: FMs can conduct in-depth analysis of legal documents and identify key information and legal concepts in the text, thereby assisting lawyers and legal advisors in more precise case analysis and legal consultation [37, 154, 155]. For example, FMs can identify clauses in contracts, extract important legal elements such as obligations, rights, and conditions, help lawyers quickly understand document contents, and identify potential legal risks. Moreover, FMs can be used for case logical reasoning, predicting possible outcomes of cases by analyzing historical cases and relevant legal provisions and providing data support for lawyers to formulate defense strategies. The ChatLaw [156] can provide real-time legal consultation and answers, helping non-professionals understand complex legal issues and even generate drafts of legal documents, reducing the workload of lawyers. Additionally, FMs can assist in legal research, quickly retrieve relevant legal literature and precedents, and provide solid evidence for legal arguments. However, LLMs can sometimes generate incorrect or fabricated information, known as “AI hallucinations”, which is a crucial challenge for domain-specific FMs in law. Another key challenge is the controversy about legal liability when using AI models for legal consultation. These challenges should be carefully addressed by practitioners in the process of developing legal models.
@@
  • Arts: The application of FMs is exploring and changing the ways of creative expression and the process of artistic production. By learning from a large number of artworks and creative concepts, FMs can generate novel artworks, music, literary works, etc., providing creative inspiration and support for artists [42]. For example, generative models can be used to produce artworks [157, 158], and fine-tuning techniques can be applied to adjust the style and content of the generated works [159, 160]. Currently, in the field of video generation, Sora, developed by OpenAI, allows users to control generated content using text, producing lifelike video works. Furthermore, FMs also show promising performance in art understanding [161, 162], which provides great potential in the field of art teaching and study. Acquiring training datasets in the art domain is highly contingent upon legal restrictions, particularly concerning copyright and intellectual property. Moreover, there is significant debate within artistic ethics about the appropriateness of employing AI models. These challenges are the key factors for FMs to contribute to art.
@@
  • Finance: FMs can cover various tasks such as risk modeling, investment strategy management, market forecasting, etc., providing powerful tools for financial institutions and investors to optimize decision-making [39, 163]. For example, FMs can be used to build credit scoring systems to assess borrowers’ credit risks. These models analyze factors such as borrowers’ historical credit records, financial conditions, and debt levels to predict their ability to repay loans and, based on this, decide whether to approve loan applications and the loan interest rates. Alternatively, FMs can be used to consider the historical performance, correlations, risks, and expected returns of various asset classes, as well as historical market data, macroeconomic indicators, political events, etc., to make optimal investor decisions. BloombergGPT, as detailed in [164], was pre-trained on a large-scale financial dataset, resulting in superior performance across many benchmarks. [165] introduces FinGPT, an open-source financial LLM, and positions it not only as a model but also as an open-source framework for Financial LLMs (FinLLMs), which has the potential to fuel innovation among researchers. Using time series analysis models to predict future stock trends and other financial information is common in the financial area. However, the accuracy of these models can significantly impact investors’ decisions. Additionally, for LLMs to serve as effective investment advisors, they must possess more robust numerical analysis capabilities, which presents a crucial challenge in leveraging their full potential.
5 Challenges ahead